How many liters of O2 at STP is needed to completely burn 20.0 L of butane(C4H10), with this reaction?


2C4H10+13O2 ---> 8CO2+10H2O


1
Expert's answer
2022-01-17T15:24:02-0500

Solution:


given chemical reaction:


"2C_4H_{10}+13O_2 \to 8CO_2 +10 H_2O"


I) We need to divide given volume of butane in order to find moles of it:


"n=\frac{20}{22.4}=0.8928moles\;of\;butane"


"n_{O_2}=\frac{0.8928 \times 13}{2}=5.799\;moles \;of\;oxygen"


II) We need to multiply moles of oxygen by volume in standard conditions


"V_{O_2}=n \times V _o=5.799 \times 22.4 L"


"V_{O_2}=129.91L"



Answer: We need 129.91 L of oxygen to completely burn 20.0 L of butane.